Please include a solution for each part

.Classify each of the frames as statically determinate, statically indeterminate, or unstable. If indeterminate, specify the d
Answer #1

For Ds = Dsi + Ose Rigid jointed plane Di = 30, Ose - frame T-3 R = 3 (number of boxes) - 301) = 3 To co ty T Be = 9-3 = 6 BY-14 TT Ds; = 3C = 18 Dse = 5-3 = 14-3=11 T To to o Degree of d static Ds - 18 +11 = 29 indeterminacy is 29 stable at the Theplease use comment section if you have any queries.